# CSGaze: Context-aware Social Gaze Prediction

### v1

@@ Abstract.
A person’s gaze offers valuable insights into their focus of attention, level of social engagement, and confidence. In this work, we investigate how contextual cues combined with visual scene and facial information can be effectively utilized to predict and interpret social gaze patterns during conversational interactions. We introduce CSGaze, a context aware multimodal approach that leverages facial, scene information as complementary inputs to enhance social gaze pattern prediction from multi-person images. The model also incorporates a fine-grained attention mechanism centered on the principal speaker, which helps in better modeling social gaze dynamics. Experimental results show that CSGaze performs competitively with state-of-the-art methods on GP-Static, UCO-LAEO and AVA-LAEO. Our findings highlight the role of contextual cues in improving social gaze prediction. Additionally, we provide initial explainability through generated attention scores, offering insights into the model’s decision-making process. We also demonstrate our model’s generalizability by testing our model on open set datasets that demonstrating its robustness across diverse scenarios. The code and data is available at https://github. com/surbhimadan92/MIP-GAF.
@@
copyright: rightsretaineddoi: 10.475/123_4isbn: 123-4567-24-567/08/06conference: 12th Indian Conference on Computer Vision, Graphics and Image Processing; December 2021; Jodhpur, Indiajournalyear: 2021price: 15.00ccs: Computing methodologies Machine learningccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
In social interactions, gaze is a crucial mechanism for expressing interest and conveying emotions. During communication, individuals continuously analyze, respond to, and observe their partner’s gaze patterns (boucher2012reach). Gaze also reflects intent—speakers often direct their gaze toward their speech target or the object they reference (lee2024towards). Several studies have explored gaze target detection, which identifies the object a person is visually focusing on within an image (so_120; so_121; so_122; so_115). However, relatively few studies have explored computational approaches for analyzing social gaze (chong2020detecting; gupta2024unified; malik2023relational).Understanding social gaze patterns has broad applications, including developing AI systems with social intelligence (gupta2024unified), assessing developmental disorders such as Autism Spectrum Disorder (ASD) (american2013diagnostic), enhancing emotion understanding (EM1; EM2), and improving engagement prediction (E1) and inferring human traits (p1; P2) in interactive settings. Psychological studies show that the frequency and duration of different gaze patterns reveal whether a conversation flows smoothly or breaks down (argyle1994gaze; kleinke1986gaze). Therefore, automatically detecting and analyzing distinct gaze behaviors is essential for enabling machines to interpret communication between individuals.
Understanding human gaze communication in social interactions is essential for several reasons: (1) it sheds light on multi-agent gaze behaviours in real-world social settings, (2) it helps robots learn human gaze patterns for more intuitive and efficient interactions, (3) it enables realistic gaze behaviour simulations in Virtual Reality, (4) it contributes to studying mental states in social contexts, and (5) it supports autism diagnosis and assessment in children (fan2019understanding). Despite its significance, social gaze prediction remains challenging due to the absence of a standardized encoding framework for gaze behaviours. Most prior studies have focused on detecting isolated gaze patterns (marin2014detecting; fan2018inferring; tafasca2023childplay), limiting their ability to generalize and identify multiple simultaneous gaze interactions in images or videos. Fan et al. (fan2019understanding) proposed six atomic-level gaze communication patterns based on cognitive psychology terminology. However, their framework lacks uniqueness, as three patterns (avert, refer, follow) represent sequential actions and can be decomposed into the remaining three (single, mutual, share), which describe static gaze states. Chang et al. (gaze) attempted to address this issue by defining inseparable gaze patterns as stationary gaze states in dyadic interactions (Shown in Figure 1), proposing that gaze behaviours in videos could be represented as sequences of frame-by-frame patterns.
@@
Social gaze prediction has traditionally relied on specific visual cues such as head orientation, scene context, and the relative positions of individuals in an image. While these approaches provide valuable insights, they often lack a comprehensive understanding of the broader contextual elements that influence gaze behaviour. A possible solution to preserve contextual cues is manual annotation or pseudo-labeling using expert models; however, these methods are resource-intensive, require task-specific models for each cue, and may not always be feasible during inference. This highlights the need for more efficient and scalable solutions. Recently, the rise of Large Language Models (LLMs) (Gaze4_33; Gaze4_34; Gaze4_12) has driven research into text-guided feature learning across various vision tasks, including crowd counting (Gaze4_28), point cloud analysis (Gaze4_56), and image generation (Gaze4_39). Despite these advancements, social gaze prediction has yet to fully harness the power of language-driven reasoning. We propose leveraging MLLMs to encode and preserve the contextual information of a scene, thereby enhancing social gaze understanding. By utilizing the zero-shot capabilities of MLLMs, we aim to incorporate linguistic context while minimizing computational overhead, paving the way for a more scalable and interpretable approach to social gaze prediction.
@@
Prompt 1.1: MLLM (Kosmos-2) Interaction System: Initialize Kosmos-2. Human: {<<prompt>> Describe how the persons are interacting in the scene.”} AI: {OUTPUT provides a textual description of the scene, detailing the objects, people, and their interactions. (See Figure 2)}
@@
Building on previous findings, this work investigates the efficacy of MLLM in context-aware social gaze prediction. We introduce CS-Gaze, a carefully designed language-vision model aimed at enhancing the semantic space of social gaze estimation. Unlike traditional approaches that focus solely on refining image encoders, our method harnesses the guidance of textual signals to improve gaze prediction. Specifically, we leverage the zero-shot capabilities of MLLMs by defining a structured prompt: “Describe how the persons are interacting in the scene.” (as shown in Prompt 1.1). This prompt extracts meaningful scene descriptions, which are then integrated into the gaze prediction framework. To further refine the prediction process, we implement a fine-grained attention mechanism that prioritizes the principal person (whose gaze is being predicted) while also considering their interaction with an associate. Additionally, a cross-modal attention-based fusion approach is employed to effectively align textual and visual representations, allowing for a richer and more contextually aware learning process. By shifting the emphasis from image encoders to language-driven contextualization, our method enhances both interpretability and performance. We evaluate CS-Gaze on the GP-Static (gaze) dataset and further test its generalizability on three additional datasets: UCO-LAEO (uco), AVA-LAEO (ava) and VSGaze (gupta2025mtgs).
@@
  • State-of-the-Art Performance: CSGaze outperforms existing methods on GP-Static dataset (gaze) and generalizes well to UCO-LAEO (uco), AVA-LAEO (ava), and VSGaze (gupta2025mtgs), demonstrating the value of language-informed context in gaze understanding.
@@ 2. Related Work
This section examines recent advancements in computer vision for identifying social gaze patterns and leveraging gaze behaviour to analyze complex social interactions.
2.1. Single Gaze Pattern
Researchers have primarily focused on two major gaze patterns in the literature: Looking-At-Each-Other (LAEO) (marin2014detecting), also known as mutual gaze, and Shared Attention (SA) (fan2018inferring). Shared attention has been studied in two ways: (1) as a classification task—determining whether shared attention is present or not—and (2) as a localization task—identifying the object or region that two or more individuals are looking at. For LAEO detection, many approaches analyze cropped head images to extract gaze direction cues. These cues are then combined with either 2D spatial data or estimated 3D geometric features to determine the LAEO label (doosti2021boosting; lin2017feature). For shared attention, several methods have integrated gaze-following heatmap predictions from all individuals (chong2020detecting; tu2022end), leading to improved performance in shared attention prediction. However, existing approaches have certain limitations, such as the inability to distinguish between multiple shared attention instances occurring simultaneously and the computational complexity of processing each person pair independently.
@@ 2.1. Single Gaze Pattern
Figure 3. CS-Gaze Framework. Phase 1: We pre-train the scene and face encoders using the GazeFollow dataset (nips15_recasens), which contains images annotated with gaze locations, providing supervisory signals for gaze estimation (See Heatmap, headlocation and gaze vector). Stage 2: Given an image with two interacting individuals: Principal (Green bounding box) and Associate (Pink bounding box), their cropped facial regions are processed through the pretrained face encoder to extract face-specific embeddings. To model the interaction between the two individuals, we employ an attention-based fusion mechanism, where the embeddings of both individuals are combined using learnable weighting parameters that determine the relative importance of each person’s facial features in the final representation.
2.2. Multiple Gaze Communication Patterns
@@
Multi-gaze analysis remains an underexplored area, with single gaze pattern analysis facing notable limitations, such as challenges in generalizing across tasks and the complexity of processing person pairs independently. These approaches often fail to capture the full spectrum of human gaze behaviour. Despite the significance of multi-gaze prediction, only a few studies have addressed this topic (fan2019understanding; gaze; gupta2025mtgs; gupta2024unified). Fan et al. (fan2019understanding) proposed a spatio-temporal graph network to hierarchically reason about gaze communication at both the atomic level (e.g., shared attention, mutual gaze) and the event level (e.g., gaze aversion, joint attention). However, their approach assumes atomic-level gaze behaviours are mutually exclusive, which is an oversimplification. For instance, person A can simultaneously maintain mutual gaze with person B while sharing attention with person C towards B.
@@
Chang et al. (gaze) redefined five static gaze patterns encompassing all possible gaze states in dyadic interactions. They introduced a network for recognizing these mutually exclusive gaze patterns from a single image and proposed the GP-Static dataset to support further research. Meanwhile, Gupta et al. (gupta2025mtgs) developed VSGaze, a dataset that provides annotations for both social gaze and gaze following, facilitating the training of more generalizable models. The social gaze labels in VSGaze include Looking-at-Head (LAH), LAEO, and SA, alongside precise gaze point annotations. Gupta et al. (gupta2024unified) introduce a unified model for gaze following and social gaze prediction by using Graph Attention Network to explicitly model people interactions, before being passed to the pairwaise task heads. To the best of our knowledge, these are the only studies focused on multi-gaze analysis. Notably, while GP-Static is publicly available, VSGaze remains inaccessible at the time of writing.
2.3. Context-Aware Social Gaze Prediction
@@
Visual gaze estimation has gained significant research interest over the past decade due to its diverse applications. While existing methods have improved prediction accuracy, they primarily rely on single-image cues to infer gaze direction, often neglecting the potential of text-based guidance, which has become increasingly relevant in vision-related tasks. To date, only two studies have explored this integration. Ruan et al. (wang2023gazeclip) proposed GazeCLIP, a text-guided gaze estimation approach leveraging the strong generalization capability of the CLIP model across vision tasks. This work was the first to distill knowledge from a large-scale vision-language model to enhance gaze estimation learning. Meanwhile, Gupta et al. (gupta2024exploring) examined the zero-shot capabilities of vision-language models (VLMs) for gaze following. Their study systematically evaluates different VLMs, prompting strategies, and in-context learning (ICL) techniques to analyze zero-shot cue recognition. They further extract contextual cues for gaze following and assess their impact when incorporated into a state-of-the-art gaze prediction model.
@@ 3.2.2. Scene Features
Building upon prior studies (chong2020detecting; gaze; so_115), which highlight the role of saliency information in gaze estimation, we introduce a scene branch to incorporate visual context from the surrounding environment. This branch processes the entire image as input and extracts scene features that complement gaze understanding, capturing crucial background elements that influence social gaze behaviour.
@@ 3.3. Proposed Method
In the first phase, we pre-train the scene and face encoders using the GazeFollow dataset (nips15_recasens), which contains images annotated with gaze locations, providing supervisory signals for gaze estimation. This pretraining step enables the encoders to learn spatially and contextually aware feature representations, capturing both the individual’s facial and head features and scene-dependent gaze cues. The model is trained to predict a $64\times 64$ heatmap representing the probability distribution of the gaze target within the image. It ensures that the encoders develop a foundational understanding of gaze behavior before being adapted for the conversational setting.
@@ 4.1. Datasets
GP-Static Dataset: The GP-Static dataset (gaze) contains 84,682 frames with 169,364 frame-level static gaze annotations, where each individual is annotated per frame. Around 20% (14,952 frames) is reserved for testing, with no overlap in TV shows or movies between training and test sets to ensure evaluation on unseen data. It defines five social gaze classes: Share, Mutual, Single, Miss, and Void (see Figure 1).
@@
VSGaze Dataset: The VSGaze dataset includes annotations for three social gaze classes: LAEO, LAH (Looking At Head) and SA. As VSGaze is not publicly available, we reconstructed it following (gupta2025mtgs), using over 800k frame samples sourced from VideoCoAtt (fan2018inferring), UCO-LAEO (uco), ChildPlay (tafasca2023childplay), and VideoTargetAttention (VAT) (chong2020detecting). As each dataset annotates only a subset of individuals, we detected missing heads using a pre-trained YOLOv5 detector and predicted gaze points with Gaze-LLE (ryan2025gaze). LAH pairs were identified when a person’s gaze intersects another’s head bounding box, LAEO pairs via mutual gaze detection, and SA pairs when both individuals’ gaze targets fall within the same head box.
@@
GF Fixed (chong2020detecting) 0.46 0.31 0.31 0.36 0.18 0.32 0.35
GF Modified (chong2020detecting) 0.61 0.26 0.29 0.42 0.34 0.38 0.43
Baseline (gaze) 0.79 0.59 0.59 0.6 0.73 0.66 0.67

### v2

@@ Abstract.
A person’s gaze offers valuable insights into their focus of attention, level of social engagement, and confidence. In this work, we investigate how contextual cues combined with visual scene and facial information can be effectively utilized to predict and interpret social gaze patterns during conversational interactions. We introduce CSGaze, a context-aware multimodal approach that leverages facial and scene information as complementary inputs to enhance social gaze pattern prediction from multi-person images. The model also incorporates a fine-grained attention mechanism centered on the principal speaker, which helps in better modeling social gaze dynamics. Experimental results show that CSGaze performs competitively with state-of-the-art methods on GP-Static, UCO-LAEO and AVA-LAEO. Our findings highlight the role of contextual cues in improving social gaze prediction. Additionally, we provide initial explainability through generated attention scores, offering insights into the model’s decision-making process. We also demonstrate our model’s generalizability by testing our model on open set datasets that demonstrate its robustness across diverse scenarios. The code and data are available at Link.
@@
journalyear: 2025copyright: ccconference: Indian Conference on Computer Vision, Graphics, and Image Processing; December 17–20, 2025; Mandi, Indiabooktitle: Indian Conference on Computer Vision, Graphics, and Image Processing (ICVGIP 2025), December 17–20, 2025, Mandi, Indiadoi: 10.1145/3774521.3774555isbn: 979-8-4007-1930-1/2025/12ccs: Computing methodologies Machine learningccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
In social interactions, gaze is a crucial mechanism for expressing interest and conveying emotions. During communication, individuals continuously analyze, respond to, and observe their partner’s gaze patterns (Boucher et al., 2012). Gaze also reflects intent—speakers often direct their gaze toward their speech target or the object they reference (Lee et al., 2024; Ghosh et al., 2022a, b). Several studies have explored gaze target detection, which identifies the object a person is visually focusing on within an image (Stiefelhagen et al., 1999, 2001; Ullah et al., 2020; Recasens et al., 2015). However, relatively few studies have explored computational approaches for analyzing social gaze (Chong et al., 2020; Gupta et al., 2024a; Malik and Isik, 2023). Understanding social gaze patterns has broad applications, including developing AI systems with social intelligence (Gupta et al., 2024a), assessing developmental disorders such as Autism Spectrum Disorder (ASD) (American Psychiatric Association et al., 2013), enhancing emotion understanding (Kumar et al., 2024a; Kataria et al., 2025), and improving engagement prediction (Kumar et al., 2024b) and inferring human traits (Kumar et al., 2024c; Kumar and Raman, 2022) in interactive settings. Psychological studies show that the frequency and duration of different gaze patterns reveal whether a conversation flows smoothly or breaks down (Argyle et al., 1994; Kleinke, 1986). Therefore, automatically detecting and analyzing distinct gaze behaviors is essential for enabling machines to interpret communication between individuals.
Understanding human gaze communication in social interactions is essential for several reasons: (1) it sheds light on multi-agent gaze behaviours in real-world social settings, (2) it helps robots learn human gaze patterns for more intuitive and efficient interactions, (3) it enables realistic gaze behaviour simulations in Virtual Reality, (4) it contributes to studying mental states in social contexts, and (5) it supports autism diagnosis and assessment in children (Fan et al., 2019). Despite its significance, social gaze prediction remains challenging due to the absence of a standardized encoding framework for gaze behaviours. Most prior studies have focused on detecting isolated gaze patterns (Marin-Jimenez et al., 2014; Fan et al., 2018; Tafasca et al., 2023), limiting their ability to generalize and identify multiple simultaneous gaze interactions in images or videos. Fan et al. (Fan et al., 2019) proposed six atomic-level gaze communication patterns based on cognitive psychology terminology. However, their framework lacks uniqueness, as three patterns (avert, refer, follow) represent sequential actions and can be decomposed into the remaining three (single, mutual, share), which describe static gaze states. Chang et al. (Chang et al., 2023) attempted to address this issue by defining inseparable gaze patterns as stationary gaze states in dyadic interactions (Shown in Figure 1), proposing that gaze behaviours in videos could be represented as sequences of frame-by-frame patterns.
@@
Social gaze prediction has traditionally relied on specific visual cues such as head orientation, scene context, and the relative positions of individuals in an image. While these approaches provide valuable insights, they often lack a comprehensive understanding of the broader contextual elements that influence gaze behaviour. A possible solution to preserve contextual cues is manual annotation or pseudo-labeling using expert models; however, these methods are resource-intensive, require task-specific models for each cue, and may not always be feasible during inference. This highlights the need for more efficient and scalable solutions. Recently, the rise of Large Language Models (LLMs) (Radford, 2018; Radford et al., 2019; Devlin, 2018) has driven research into text-guided feature learning across various vision tasks, including crowd counting (Liang et al., 2023), point cloud analysis (Zhang et al., 2022), and image generation (Ramesh et al., 2022). Despite these advancements, social gaze prediction has yet to fully harness the power of language-driven reasoning. We propose leveraging MLLMs to encode and preserve the contextual information of a scene, thereby enhancing social gaze understanding. By utilizing the zero-shot capabilities of MLLMs, we aim to incorporate linguistic context while minimizing computational overhead, paving the way for a more scalable and interpretable approach to social gaze prediction.
Prompt 1.1: MLLM (Kosmos-2) Interaction
System: Initialize Kosmos-2.
Human: {<<prompt>> Describe how the persons are interacting in the scene.”}
@@
AI: {OUTPUT provides a textual description of the scene, detailing the objects, people, and their interactions. (See Figure 2)}
@@
Building on previous findings, this work investigates the efficacy of MLLM in context-aware social gaze prediction. We introduce CS-Gaze, a carefully designed language-vision model aimed at enhancing the semantic space of social gaze estimation. Unlike traditional approaches that focus solely on refining image encoders, our method harnesses the guidance of textual signals to improve gaze prediction. Specifically, we leverage the zero-shot capabilities of MLLMs by defining a structured prompt: “Describe how the persons are interacting in the scene.” (as shown in Prompt 1.1). This prompt extracts meaningful scene descriptions, which are then integrated into the gaze prediction framework. To further refine the prediction process, we implement a fine-grained attention mechanism that prioritizes the principal person (whose gaze is being predicted) while also considering their interaction with an associate. Additionally, a cross-modal attention-based fusion approach is employed to effectively align textual and visual representations, allowing for a richer and more contextually aware learning process. By shifting the emphasis from image encoders to language-driven contextualization, our method enhances both interpretability and performance. We evaluate CS-Gaze on the GP-Static (Chang et al., 2023) dataset and further test its generalizability on three additional datasets: UCO-LAEO (Marin-Jimenez et al., 2019a), AVA-LAEO (Marin-Jimenez et al., 2019b) and VSGaze (Gupta et al., 2025).
@@
  • State-of-the-Art Performance: CSGaze outperforms existing methods on GP-Static dataset (Chang et al., 2023) and generalizes well to UCO-LAEO (Marin-Jimenez et al., 2019a), AVA-LAEO (Marin-Jimenez et al., 2019b), and VSGaze (Gupta et al., 2025), demonstrating the value of language-informed context in gaze understanding.
@@ 2. Related Work
This section examines recent advancements (Ghosh et al., 2023) in computer vision for identifying social gaze patterns and leveraging gaze behaviour to analyze complex social interactions.
Single Gaze Pattern. Researchers have primarily focused on two major gaze patterns in the literature: Looking-At-Each-Other (LAEO) (Marin-Jimenez et al., 2014), also known as mutual gaze, and Shared Attention (SA) (Fan et al., 2018). Shared attention has been studied in two ways: (1) as a classification task—determining whether shared attention is present or not—and (2) as a localization task—identifying the object or region that two or more individuals are looking at. For LAEO detection, many approaches analyze cropped head images to extract gaze direction cues. These cues are then combined with either 2D spatial data or estimated 3D geometric features to determine the LAEO label (Doosti et al., 2021; Lin et al., 2017). For shared attention, several methods have integrated gaze-following heatmap predictions from all individuals (Chong et al., 2020; Tu et al., 2022), leading to improved performance in shared attention prediction. However, existing approaches have certain limitations, such as the inability to distinguish between multiple shared attention instances occurring simultaneously and the computational complexity of processing each person pair independently.
@@
Figure 3. CS-Gaze Framework. Phase 1: We pre-train the scene and face encoders using the GazeFollow dataset (Recasens et al., 2015), which contains images annotated with gaze locations, providing supervisory signals for gaze estimation (See Heatmap, headlocation and gaze vector). Stage 2: Given an image with two interacting individuals: Principal (Green bounding box) and Associate (Pink bounding box), their cropped facial regions are processed through the pretrained face encoder to extract face-specific embeddings. To model the interaction between the two individuals, we employ an attention-based fusion mechanism, where the embeddings of both individuals are combined using learnable weighting parameters that determine the relative importance of each person’s facial features in the final representation.
@@
Multiple Gaze Communication Patterns. Multi-gaze analysis remains an underexplored area, with single gaze pattern analysis facing notable limitations, such as challenges in generalizing across tasks and the complexity of processing person pairs independently. These approaches often fail to capture the full spectrum of human gaze behaviour. Despite the significance of multi-gaze prediction, only a few studies have addressed this topic (Fan et al., 2019; Chang et al., 2023; Gupta et al., 2025, 2024a). Fan et al. (Fan et al., 2019) proposed a spatio-temporal graph network to hierarchically reason about gaze communication at both the atomic level (e.g., shared attention, mutual gaze) and the event level (e.g., gaze aversion, joint attention). However, their approach assumes atomic-level gaze behaviours are mutually exclusive, which is an oversimplification. For instance, person A can simultaneously maintain mutual gaze with person B while sharing attention with person C towards B.
@@
Chang et al. (Chang et al., 2023) redefined five static gaze patterns encompassing all possible gaze states in dyadic interactions. They introduced a network for recognizing these mutually exclusive gaze patterns from a single image and proposed the GP-Static dataset to support further research. Meanwhile, Gupta et al. (Gupta et al., 2025) developed VSGaze, a dataset that provides annotations for both social gaze and gaze following, facilitating the training of more generalizable models. The social gaze labels in VSGaze include Looking-at-Head (LAH), LAEO, and SA, alongside precise gaze point annotations. Gupta et al. (Gupta et al., 2024a) introduce a unified model for gaze following and social gaze prediction by using Graph Attention Network to explicitly model people interactions, before being passed to the pairwaise task heads. To the best of our knowledge, these are the only studies focused on multi-gaze analysis. Notably, while GP-Static is publicly available, VSGaze remains inaccessible at the time of writing.
@@
Context-Aware Social Gaze Prediction. Visual gaze estimation has gained significant research interest over the past decade due to its diverse applications. While existing methods have improved prediction accuracy, they primarily rely on single-image cues to infer gaze direction, often neglecting the potential of text-based guidance, which has become increasingly relevant in vision-related tasks. To date, only two studies have explored this integration. Ruan et al. (Wang et al., 2023) proposed GazeCLIP, a text-guided gaze estimation approach leveraging the strong generalization capability of the CLIP model across vision tasks. This work was the first to distill knowledge from a large-scale vision-language model to enhance gaze estimation learning. Meanwhile, Gupta et al. (Gupta et al., 2024b) examined the zero-shot capabilities of vision-language models (VLMs) for gaze following. Their study systematically evaluates different VLMs, prompting strategies, and in-context learning (ICL) techniques to analyze zero-shot cue recognition. They further extract contextual cues for gaze following and assess their impact when incorporated into a state-of-the-art gaze prediction model.
@@ 3.2.2. Scene Features
Building upon prior studies (Chong et al., 2020; Chang et al., 2023; Recasens et al., 2015), which highlight the role of saliency information in gaze estimation, we introduce a scene branch to incorporate visual context from the surrounding environment. This branch processes the entire image as input and extracts scene features that complement gaze understanding, capturing crucial background elements that influence social gaze behaviour.
@@ 3.3. Proposed Method
In the first phase, we pre-train the scene and face encoders using the GazeFollow dataset (Recasens et al., 2015), which contains images annotated with gaze locations, providing supervisory signals for gaze estimation. This pretraining step enables the encoders to learn spatially and contextually aware feature representations, capturing both the individual’s facial and head features and scene-dependent gaze cues. The model is trained to predict a $64\times 64$ heatmap representing the probability distribution of the gaze target within the image. It ensures that the encoders develop a foundational understanding of gaze behavior before being adapted for the conversational setting.
@@ 4.1. Datasets
GP-Static Dataset: The GP-Static dataset (Chang et al., 2023) contains 84,682 frames with 169,364 frame-level static gaze annotations, where each individual is annotated per frame. Around 20% (14,952 frames) is reserved for testing, with no overlap in TV shows or movies between training and test sets to ensure evaluation on unseen data. It defines five social gaze classes: Share, Mutual, Single, Miss, and Void (see Figure 1).
@@
VSGaze Dataset: The VSGaze dataset includes annotations for three social gaze classes: LAEO, LAH (Looking At Head) and SA. As VSGaze is not publicly available, we reconstructed it following (Gupta et al., 2025), using over 800k frame samples sourced from VideoCoAtt (Fan et al., 2018), UCO-LAEO (Marin-Jimenez et al., 2019a), ChildPlay (Tafasca et al., 2023), and VideoTargetAttention (VAT) (Chong et al., 2020). As each dataset annotates only a subset of individuals, we detected missing heads using a pre-trained YOLOv5 detector and predicted gaze points with Gaze-LLE (Ryan et al., 2025). LAH pairs were identified when a person’s gaze intersects another’s head bounding box, LAEO pairs via mutual gaze detection, and SA pairs when both individuals’ gaze targets fall within the same head box.
@@
GF Fixed (Chong et al., 2020) 0.46 0.31 0.31 0.36 0.18 0.32 0.35
GF Modified (Chong et al., 2020) 0.61 0.26 0.29 0.42 0.34 0.38 0.43
Baseline (Chang et al., 2023) 0.79 0.59 0.59 0.6 0.73 0.66 0.67